# ConDiSR: Contrastive Disentanglement and Style Regularization for Single Domain Generalization

### v1

@@ 3 Experimental Setup
Vanilla ERM $83.1_{\pm 0.9}$ $65.2_{\pm 1.1}$ $58.0_{\pm 0.4}$ $64.8_{\pm 1.9}$ $56.2_{\pm 0.8}$ $65.4$
ERM $87.0_{\pm 0.7}$ $81.0_{\pm 2.2}$ $89.3_{\pm 0.4}$ $74.1_{\pm 1.5}$ $88.8_{\pm 0.5}$ $84.1$
$\text{C}^{2}\text{SDG}$ $\mathbf{90.2_{\pm 1.3}}$ $77.2_{\pm 1.0}$ $\mathbf{90.5_{\pm 0.3}}$ $72.5_{\pm 2.8}$ $89.4_{\pm 0.3}$ $83.9$
ConDiSR $87.4_{\pm 0.6}$ $\mathbf{81.8_{\pm 1.0}}$ $90.4_{\pm 0.7}$ $\mathbf{75.0_{\pm 0.8}}$ $\mathbf{89.7_{\pm 0.6}}$ $\mathbf{84.9}$
4 Results and Discussion
Table 1 compares performance of our method versus ERM, $\text{C}^{2}$ [5] and ERM without augmentations (Vanilla ERM) in SDG setting for each separate center being set as the source domain. Each value is achieved by averaging three experimental results with different seeds.
As it can be seen from the Table 1, our method consistently outperforms baselines by the margin of $\sim 1\%$ as well as shows more stable performance with less deviation across the runs. We argue that reconstruction-based regularization aids the disentanglement module of the framework by enforcing stronger dependency between the input images and style-dependent channels of the shallow-level feature map. Due to direct interconnection this improves extraction of the structure-related channels as well.
The qualitative comparison of performance between $\text{C}^{2}\text{SDG}$ and ConDiSR on the Figure 3 shows that our method pays more attention to areas, containing semantically important information, which allows us to claim that the style regularization does indeed improve extraction of structural features as well.
The comparison between Vanilla ERM, ERM and $\text{C}^{2}\text{SDG}$ shows that the latter two mehtods achieve a significantly better performance than Vanilla ERM. However, when compared to each other, ERM and $\text{C}^{2}\text{SDG}$ have almost the same average accuracy, which leads us to believe that when it comes to the binary classification task, the primary driver of performance improvement of $\text{C}^{2}\text{SDG}$ might be extensive augmentations as well as increased number of training iterations resulting from the augmentation strategy in [5] rather than the idea of channel-wise contrastive disentanglement in its default version.
Table 2: Evaluation of the effect of changing the reconstruction resolution in ConDiSR. Standard deviation over three seeds is given in the subscript. The best accuracy (%) is highlighted in bold.
@@ 4 Results and Discussion
Table 3: Investigation of the effect of combining various style augmentation techniques with ConDiSR. Standard deviation over three seeds is given in the subscript. The best accuracy (%) is highlighted in bold.
@@
+ ConDiSR $\mathbf{87.2_{\pm 0.1}}$ $\mathbf{82.8_{\pm 0.5}}$ $\mathbf{92.1_{\pm 0.6}}$ $\mathbf{75.6_{\pm 0.6}}$ $90.6_{\pm 0.5}$ $\mathbf{85.7}$
@@
+ ConDiSR $84.5_{\pm 0.7}$ $\mathbf{84.2_{\pm 0.7}}$ $90.7_{\pm 0.4}$ $75.8_{\pm 0.8}$ $89.5_{\pm 0.6}$ $\mathbf{84.9}$
@@
+ ConDiSR $\mathbf{86.0_{\pm 0.4}}$ $\mathbf{85.7_{\pm 0.5}}$ $\mathbf{90.8_{\pm 0.7}}$ $\mathbf{78.1_{\pm 0.6}}$ $87.7_{\pm 0.9}$ $\mathbf{85.7}$
4.0.1 Reconstruction resolution.
From the Table 2 we can see that the optimal reconstruction resolution for our method is 48, which supports our preliminary assumption, that while performing the reconstruction-based style regularization, lack of any downsampling makes the style-related feature maps learn unnecessary structural features and hampers the work of disentanglement module. Conversely, if the target resolution is too low, the downsampling removes too much information from the source image, thereby diminishing the effect of style regularization technique.
4.0.2 Style augmentation.
@@
As it can be seen in the Table 3, application of various style-augmentation methods [24] to the feature maps after each of the first three layers of the network does not provide any significant improvement compared to the baselines. However in conjunction with ConDiSR, it results in the highest accuracy reached, which raises an assumption that despite the usage of the disentanglement technique, structure-related feature maps extracted by our method still carry various important style-dependent features that are correlated with the classification label and benefit from augmentations.
[…]
@@ 5 Conclusion
This paper presents a novel method, ConDiSR, designed for single domain generalization in the classification of images. Leveraging contrastive channel-wise disentanglement alongside style regularization through low-resolution reconstruction, ConDiSR demonstrates an improvement over baseline methods, achieving an increase in accuracy of approximately 1%. Successful application of ConDiSR shows the potential of disentanglement-based methodologies in enhancing model performance and generalizability, particularly in classification tasks within complex and heterogeneous datasets such as histopathology images. Our results underscore the importance of further exploration of disentanglement-based approaches and encourage further research in this domain. Potential future work may involve evaluation of the proposed method’s performance on other classification datasets in order to test its generalizability.

### v3

@@ 4 Results and Discussion
ERM $96.5_{\pm 0.2}$ $89.1_{\pm 1.2}$ $94.8_{\pm 0.5}$ $92.8_{\pm 0.4}$ $78.7_{\pm 2.3}$ $90.4$
$\text{C}^{2}\text{SDG}$ $96.3_{\pm 0.5}$ $89.4_{\pm 2.0}$ $94.3_{\pm 1.0}$ $90.3_{\pm 2.6}$ $79.7_{\pm 2.6}$ $90.0$
Ours $\mathbf{97.3_{\pm 0.5}}$ $\mathbf{92.1_{\pm 0.7}}$ $\mathbf{96.1_{\pm 0.3}}$ $\mathbf{94.2_{\pm 0.3}}$ $\mathbf{84.1_{\pm 0.3}}$ $\mathbf{92.8}$
Table 3: Performance comparison in MSDG setting on Camelyon17-WILDS. Each column represents accuracy on domain $C_{i}$ after training on the other 4 domains
Aptos EyePACS Messidor Messidor2 Average
ERM $64.4_{\pm 1.4}$ $50.1_{\pm 0.9}$ $62.7_{\pm 1.8}$ $62.2_{\pm 0.5}$ $59.9$
$\text{C}^{2}\text{SDG}$ $60.2_{\pm 2.3}$ $48.5_{\pm 1.8}$ $61.7_{\pm 2.5}$ $60.3_{\pm 1.6}$ $57.7$
Ours $\mathbf{66.1_{\pm 1.0}}$ $\mathbf{51.8_{\pm 0.4}}$ $\mathbf{63.0_{\pm 0.8}}$ $\mathbf{63.3_{\pm 0.8}}$ $\mathbf{61.1}$
Table 4: Performance comparison in MSDG setting on the composition of DR datasets. Each column represents accuracy on the respective domain after training on the other 3 domains
Table 5: Evaluation of the effect of changing the reconstruction resolution in ConDiSR is SDG setting. Each column represents performance after training on the domain $C_{i}$ averaged across the remaining 4 domains.
@@
Table 6: Investigation of the effect of combining various style augmentation techniques with ConDiSR. Each column represents performance after training on the domain $C_{i}$ averaged across the remaining 4 domains.
@@
+ Ours $\mathbf{87.2_{\pm 0.1}}$ $\mathbf{82.8_{\pm 0.5}}$ $\mathbf{92.1_{\pm 0.6}}$ $\mathbf{75.6_{\pm 0.6}}$ $90.6_{\pm 0.5}$ $\mathbf{85.7}$
@@
+ Ours $84.5_{\pm 0.7}$ $\mathbf{84.2_{\pm 0.7}}$ $90.7_{\pm 0.4}$ $75.8_{\pm 0.8}$ $89.5_{\pm 0.6}$ $\mathbf{84.9}$
@@
+ Ours $\mathbf{86.0_{\pm 0.4}}$ $\mathbf{85.7_{\pm 0.5}}$ $\mathbf{90.8_{\pm 0.7}}$ $\mathbf{78.1_{\pm 0.6}}$ $87.7_{\pm 0.9}$ $\mathbf{85.7}$
Table 1 compares performance of our method versus ERM and $\text{C}^{2}$ [8] in SDG setting for each separate medical center being set as the source domain. Each value is achieved by averaging three experimental results with different seeds.
As it can be seen from the tables, our method consistently outperforms baselines by 1-2% on average. Moreover it shows more stable performance with significantly smaller deviation across the runs. We argue that reconstruction-based regularization aids the disentanglement module of the framework by enforcing stronger dependency between the input images and style-dependent channels of the shallow-level feature map. Due to direct interconnection this improves extraction of the structure-related channels as well.
Figure 5: Qualitative performance comparison between ConDiSR and C2SDG [8] via Grad-CAM [16] with the tumor presense masks taken from the original Camelyon17 challenge dataset [1].
The comparison between ERM and $\text{C}^{2}\text{SDG}$ shows that both mehtods achieve very close performance scores, which leads us to believe that when it comes to the binary classification task, the primary driver of good performance of $\text{C}^{2}\text{SDG}$ might be extensive augmentations as well as increased number of training iterations resulting from the augmentation strategy in [8] rather than the idea of channel-wise contrastive disentanglement in its default version.
The qualitative comparison of performance between $\text{C}^{2}\text{SDG}$ and ConDiSR is shown on the Figure 5. It is obtained by visualizing gradient activation maps [16] on samples with positive tumor label. Comparison with segmentation masks from original Camelyon17 challenge dataset shows that our method pays more attention to areas, containing semantically important information.
Table 2 compares performance of our method versus ERM and $\text{C}^{2}$ in SDG setting on the combination of the Diabetic Retinopathy datasets. Each of the datasets is one-by-one set as the source domain. As it can be seen from the tables, our method consistently outperforms the baselines while showing higher stability with lower deviation in results. The improvement is especially noticeable with Messidor2 as the source domain.
Multi-Source Domain Generlaization. The results of the experiments in MSDG setting are shown in the Tables 3 and 4. As it can be seen from the Table 3 when applied to Camelyon17-WILDS our method noticeably outperforms baselines on each of the tagret domains. The difference is epecially noticeable on the data from the medical center 4, which overall is the hardest to generalize to for all of the methods. The average accuracy improvement is from $90.4\%$ to $92.8\%$ which corresponds to approximately $20\%$ classification error reduction. When it comes to DR grading, our method again outperforms the baselines by the margin of more than $1\%$.
Reconstruction resolution. From the Table 5 we can see that the optimal reconstruction resolution for our method is 48x48. This supports our preliminary assumption, that reconstruction of the input images of the original size (96x96) forces unnecessary structural features into $f_{sty}$ and hampers the work of disentanglement module. On the other hand, reconstruction of the images resized to the resolutions that are too small (24x24) does not help the model’s performance. Similarly to [6], the gap seems to arise between the tasks of recognition and reconstruction. When the target image resolution is too low, the decoder of the reconstruction model is too shallow (as it consists of fewer upsampling blocks) which makes the input of the reconstruction network, $f_{sty}$, too specialized for reconstruction but not particularly informative from the point of view of recognition.
It is important to clarify, that Table 5 shows performance in the SDG setting with each column representing the source domain and the accuracy being computed over the rest of the domains unlike it is presented in the Tables 1 and 2.
@@
Style augmentation. As it can be seen in the Table 6, application of various style-augmentation methods [28] to the feature maps after each of the first three layers of the network does not provide any significant improvement compared to the baselines. However in conjunction with ConDiSR, it results in the highest accuracy reached, which raises an assumption that despite the usage of the disentanglement technique, structure-related feature maps extracted by our method still carry various important style-dependent features that are correlated with the classification label and benefit from augmentations. Similarly to Table 5, Table 6 shows performance in the SDG setting with each column representing the source domain and the accuracy being computed over the rest of the domains.
@@ 5 Conclusion
This paper presents a novel method, ConDiSR, designed for single domain generalization in the classification of images. Leveraging contrastive channel-wise disentanglement alongside style regularization through low-resolution reconstruction, ConDiSR demonstrates an improvement over baseline methods, achieving an increase in accuracy in both SDG and MSDG settings on various medical imaging modalitites. Successful application of ConDiSR shows the potential of disentanglement-based methodologies in enhancing model performance and generalizability, particularly in classification tasks within complex and heterogeneous datasets such as histopathology images. Our results underscore the importance of further exploration of disentanglement-based approaches and encourage further research in this domain. Potential future work may involve evaluation of the proposed method’s performance on other classification datasets in order to test its generalizability.